# Attributing Response to Context: A Jensen–Shannon Divergence Driven Mechanistic Study of Context Attribution in Retrieval-Augmented Generation

### v2

@@ 6.2 Mutually Verifying JSD-based Mechanistic Study via the Semantic Gains of Attention and MLPs
Finally, we can calculate Spearman $\rho$ of $J_{\text{Top}-N}(\text{MLP})\cap S_{\text{Top}-N}^{(+)}$ and $G_{\text{Top}-N}(\text{MLP})\cap S_{\text{Top}-N}^{(+)}$, where $S_{\text{Top}-N}^{(+)}=\text{sort}(S^{(+)},\text{ascending})$. For attention components, we first average the JSD scores of all attention heads in the same layer to build $\{\text{JSD}_{\text{Attn}}^{\ell}\}_{l=0}^{L}=\{\frac{1}{H}\sum_{h=0}^{H}\text{JSD}_{\text{Attn}}^{\ell,h}\}_{l=0}^{L}$, and then further calculate Spearman $\rho$ of $J_{\text{Top}-N}(\text{Attn})\cap S_{\text{Top}-N}^{(+)}$ and $G_{\text{Top}-N}(\text{Attn})\cap S_{\text{Top}-N}^{(+)}$.
The benefit of using consensus $S^{(+)}$ instead of the raw JSD or semantic gain values is that $S^{(+)}$ will remove all scaling issue due to the different units and variances of JSD or semantic gains, and a single extremely large JSD or semantic gain will not swamp the fusion, which is robust to outliers. Table 3 shows that the $\rho$ of overlap of top-10 located attention and MLP layers about JSD-based and semantic gain-based metrics with consensus are statistically significant or highly significant, which further verifies the effectiveness of our proposed JSD-based mechanistic approach.
Table 3: Spearman’s $\rho$ of the overlap about top-10 located attentions and MLPs between JSD-based mechanistic and semantic gain-based metrics over all datasets and RAG models. $\diamondsuit$ and $\spadesuit$ indicate $p$-value is $<$ 0.05 and $<$ 0.01, respectively.
Modules Top-$10$ Datasets Qwen2 1.5B IT Qwen2 7B IT Gemma2 2B IT Gemma2 9B IT
Attention TyDi QA 6.83$\diamondsuit$ 7.26$\diamondsuit$ 6.91$\diamondsuit$ 7.31$\diamondsuit$ 7.62$\spadesuit$ 7.25$\diamondsuit$ 7.63$\spadesuit$ 7.28$\diamondsuit$
Hotpot QA 6.73$\diamondsuit$ 6.65$\diamondsuit$ 6.81$\diamondsuit$ 6.79$\diamondsuit$ 6.68$\diamondsuit$ 6.67$\diamondsuit$ 6.72$\diamondsuit$ 6.73$\diamondsuit$
MuSiQue 6.67$\diamondsuit$ 6.72$\diamondsuit$ 6.72$\diamondsuit$ 6.83$\diamondsuit$ 6.69$\diamondsuit$ 6.71$\diamondsuit$ 6.73$\diamondsuit$ 6.75$\diamondsuit$
MLP TyDi QA 6.90$\diamondsuit$ 7.72$\spadesuit$ 6.96$\diamondsuit$ 7.67$\spadesuit$ 7.75$\spadesuit$ 8.03$\spadesuit$ 7.78$\spadesuit$ 8.05$\spadesuit$
Hotpot QA 6.83$\diamondsuit$ 7.49$\spadesuit$ 6.87$\diamondsuit$ 7.52$\spadesuit$ 7.50$\spadesuit$ 8.02$\spadesuit$ 7.53$\spadesuit$ 8.06$\spadesuit$
MuSiQue 6.87$\diamondsuit$ 7.12$\diamondsuit$ 6.91$\diamondsuit$ 7.18$\diamondsuit$ 7.51$\spadesuit$ 8.04$\spadesuit$ 7.54$\spadesuit$ 8.05$\spadesuit$
7 Case Studies of Located Attention Heads and MLPs for Context Attribution
Based on the semantic gains analysis from the § 6.2, we further visualise the projection of middle-residual stream $\mathbf{x}_{i}^{\ell,\text{mid}}$ and post-residual stream $\mathbf{x}_{i}^{\ell,\text{post}}$ via Logit Lens to vocabulary space in the Figure 4 and Appendix J. In the Figure 4, Qwen2-1.5B-IT was given a data from TyDi QA dev dataset with the context about mosquitos introduction from Wikipedia and query “How many wings does a mosquito have?” as input, and it generates responses “A mosquito has two wings.” as output. Based on our proposed ARC-JSD method, we successfully located the top-relevant context sentence, i.e., “Mosquitoes have a slender segmented body, a pair of wings, three pairs of long hair-like legs, feathery antennae, and elongated mouthparts”. When we compare the heatmap between $\mathbf{x}_{i}^{\ell,\text{post}}$ and $\mathbf{x}_{i}^{\ell,\text{mid}}$ in Figure 4 from Layer 20 to Layer 27 (See Appendix J for the whole heatmap), we can find that the probability of correct token is increased significantly after the $\mathbf{x}_{i}^{\ell,\text{post}}$ compared to $\mathbf{x}_{i}^{\ell,\text{mid}}$, such as ‘wings’ in Layer 23, ‘A’, ‘has’, ‘two’ in Layer 26, and ‘mosquito’, ‘two’, ‘A’ in Layer 27, which aligns with our findings that MLP contribute more for context attribution in higher layers using JSD-based metric from the § 5.2. In addition, we can find that several correct tokens are gradually transferred from their Chinese format to the English version in Qwen2 models, such as ‘{CJK*}UTF8gbsn一只 (A)’, ‘{CJK*}UTF8gbsn拥有 (has)’ and ‘{CJK*}UTF8gbsn翅膀 (wings)’, which is reasonable as Chinese is one of main language resources used in the Qwen2 model pre- and post-training (yang2024qwen2, ). This finding also matches observations from Wu et al. wu2025the that representations tend to be anchored by semantically-equivalent dominant-language tokens in higher layers. Moreover, we conduct an ablation study to compare the JSD difference of responses by masking the top-10 relevant attention heads and randomly-selected 10 attention heads. Generally, attention heads using JSD-based metric cause larger JSD scores compared to the random attention heads, which further verifies the effectiveness of our proposed ARC-JSD method (see Appendix I for details).
Figure 4: The projection of middle-residual stream $\mathbf{x}_{i}^{\ell,\text{mid}}$ and post-residual stream $\mathbf{x}_{i}^{\ell,\text{post}}$ via Logit Lens to vocabulary space from layer 20 to layer 27 of Qwen2-1.5B IT in TyDi QA data sample, where the generated response $\mathcal{R}$ is “A mosquito has two wings.” (See Appendix J for all layer projections). Each cell shows the most probable token decoded via Logit Lens. The colour indicates the probability of the decoded token of the corresponding $\mathbf{x}_{i}^{\ell,\text{mid}}$ or $\mathbf{x}_{i}^{\ell,\text{post}}$ via Logit Lens.
@@ 8 Conclusion
This study introduces ARC-JSD, an inference-time JSD–based metric that attributes responses in RAG directly to their context sentences without additional fine-tuning or surrogate modelling. Evaluations on diverse QA benchmarks and multiple scales of instruction-tuned LLMs demonstrate that ARC-JSD achieves higher attribution accuracy while markedly reducing computational overhead relative to surrogate approaches. Combined with the Logit Lens, ARC-JSD further isolates relevant attention heads and MLPs underpinning context attribution, thereby advancing mechanistic interpretability. Collectively, these findings enhance the transparency of RAG systems and lay groundwork for future research on reliable, efficient RAG models.
@@ Acknowledgement
Appendix A Broad Impact
RAG systems underpin a wide range of everyday activities, from itinerary planning and news aggregation to document drafting, by combining LLMs reasoning with evidence retrieved from external sources. Yet, the practical value of these systems hinges on our ability to verify that each generated statement is genuinely grounded in the retrieved material. The proposed post-hoc ARC-JSD method offers a lightweight, modular solution to this problem. Because ARC-JSD can be seamlessly integrated into any open-source RAG pipeline, it provides developers and researchers with an immediate way of auditing attribution fidelity, thereby strengthening the transparency, reliability, and ultimately the public trust in RAG-based applications.
Appendix B Limitations
Our work focuses on the analysis to (i) identify the context sentences that most strongly influence a RAG model’s output and (ii) attribute that influence to specific attention heads and MLP layers via a JSD-based metric. Two important directions, therefore, remain unexplored. First, our layer-level view does not reveal which individual neurons within the MLPs mediate context attribution; techniques such as sparse autoencoder (SAE) probing could provide the necessary resolution. Second, we have not yet examined whether surgical interventions on the identified attention heads, or on the putative neuron-level circuits, can be used to steer or constrain the model’s behaviour. Addressing these questions would deliver a more fine-grained mechanistic understanding and open the door to reliable, attribution-aware editing of RAG systems.
Appendix C Details of the Internal Mechanisms of LLMs
We consider the standard autoregressive Transformer architecture used in LLMs, originally introduced by [27] and subsequently analysed in a series of mechanistic studies [11, 7, 10, 6, 15, 16, 39]. Given a prompt of length $T$, the input tokens $(t_{1},\dots,t_{T})$ from the context-query pair, each drawn from a vocabulary $\mathcal{V}$, are mapped to $d$-dimensional embedding vectors $\mathbf{x}^{0}_{i}\in\mathbb{R}^{d}$, where the embedding matrix $W_{E}\in\mathbb{R}^{|\mathcal{V}|\times d}$.
@@ Appendix C Details of the Internal Mechanisms of LLMs
where $\mathbf{a}^{\ell}_{i}$ and $\mathbf{m}^{\ell}_{i}$ denote the contributions of the attention and MLP sub-modules, respectively. [^3] Layer normalisation preceding each sub-module is omitted here for clarity.
After the final layer, a LayerNorm $\sigma(\cdot)$ and the unembedding matrix $W_{U}\in\mathbb{R}^{d\times|\mathcal{V}|}$ produce the next-token distribution
@@
A concise per-head summary is $\mathbf{a}^{\ell,h}_{i}\;=\;\sum_{j}\mathbf{a}^{\ell,h}_{i,j}$.
@@
Appendix D Experimental Details
@@
Appendix E GPT-4.1 as Judge for Comparison between Generated Responses of RAG models and Gold Answers from Datasets
@@
Appendix F Computational Efficiency Between Contextcite and Our ARC-JSD
We mainly compare the computational efficiency between the Contextcite [5] and our proposed ARC-JSD when attributing responses to relevant context. As Figure 5 shows, our ARC-JSD method can achieve up to 3-fold speedup compared to the Contextcite baseline. The main reason is that our ARC-JSD only needs $|\mathcal{C}|+1$ RAG forward calls to locate top-relevant context, where $|\mathcal{C}|$ is significantly smaller than Contextcite’s $n$ calls ($n+1=256$ can achieve a stable RMSE based on their work [5]).
Figure 5: Computational efficiency comparison between our ARC-JSD and Contextcite ($n=256$ calls) across all datasets and RAG models. Each bar represents the average seconds per sample for ARC-JSD and Contextcite with different RAG models and datasets.
Appendix G Examples of ARC-JSD Context Attribution
@@
… The word “mosquito” (formed by mosca and diminutive -ito)[2] is Spanish for “little fly”. [3] Mosquitoes have a slender segmented body, a pair of wings, three pairs of long hair-like legs, feathery antennae, and elongated mouthparts.
@@
… His mother’s name was not recorded in history. Title: Mauro Scocco Mauro Scocco (born 11 September 1962) is a Swedish pop artist of Italian descent. He has been described as “one of the sharpest songwriters in Sweden”. …
@@
… It was written by Bieber, Julia Michaels, Justin Tranter and BloodPop, with production handled by BloodPop. The song was released through GENPOP Corporation, RBMG Records, School Boy Records, Def Jam Recordings and Republic Records on August 17, 2017. It is Bieber’s first single credited as lead artist to be released in over a year, the previous being "Company", which was released in March 2016 for his studio album "Purpose" (2015). …
@@
… It is also known in Brazil as "Chapolin", "Vermelhinho" ("Little Red One") and "Polegar Vermelho" ("Red Thumb") in allusion to the famous fairy tale character Tom Thumb. Title: Albertsons Albertsons Companies LLC is an American grocery company founded and based in Boise, Idaho. It is privately owned and operated by investors, including Cerberus Capital Management.
Appendix H JSD-based Mechanistic Insights for Located Attention Heads and MLPs
@@
Appendix I JSD Comparison between Masking Located Attention Heads and Random Attention Heads
We conducted an ablation study to compare the JSD difference by masking the top-10 relevant attention heads and randomly-selected 10 attention heads. Results show that top-10 attention heads located by the JSD-based metric have higher JSD scores of the same responses while masking in the Table 5.
Masking Top-10 Relevant Attention Heads Randomly Masking 10 Attention Heads
2.23$\pm$0.12 1.53$\pm$0.76
@@ Appendix I JSD Comparison between Masking Located Attention Heads and Random Attention Heads
Appendix J Case Studies of Attention and MLP’s Contribution for Each Response Token
@@
Figure 12: The projection of middle-residual stream $\mathbf{x}_{i}^{\ell,\text{mid}}$ and post-residual stream $\mathbf{x}_{i}^{\ell,\text{post}}$ via Logit Lens to vocabulary space from layer 20 to layer 27 of Qwen2-1.5B IT in TyDi QA data sample, where the generated response $\mathcal{R}$ is “A mosquito has two wings.”. Each cell shows the most probable token decoded via Logit Lens. The colour indicates the probability of the decoded token of the corresponding $\mathbf{x}_{i}^{\ell,\text{mid}}$ or $\mathbf{x}_{i}^{\ell,\text{post}}$ via Logit Lens.
@@
Figure 14: The projection of middle-residual stream $\mathbf{x}_{i}^{\ell,\text{mid}}$ and post-residual stream $\mathbf{x}_{i}^{\ell,\text{post}}$ via Logit Lens to vocabulary space from layer 20 to layer 27 of Qwen2-1.5B IT in TyDi QA data sample, where the generated response $\mathcal{R}$ is “The Meiji Restoration took place in Japan.”. Each cell shows the most probable token decoded via Logit Lens.

### v4

@@ 7 Case Studies of Located Attention Heads and MLPs
To further evaluate how the located internal components affect RAG behaviours, we conduct extra experiments on Qwen2 7B IT to demonstrate whether ARC-JSD can reduce hallucinations and maintain factual information at the same time. We select 200 data samples from Hotpot QA where Qwen2 7B IT did not provide a truthful response compared to the ground truth. From Table 4, we can use the JSD attribution scores and semantic gain $G$ as a confidence gate to reduce the hallucination rate by $\approx 39\%$ without hurting factual F1 compared to random gating attention and MLPs.  [^4] This experiment mainly evaluates how our ARC-JSD located attention and MLP affect the hallucinations and factual information in the RAG response based on the well-known agreement that attention heads and MLP focus on retrieval behavior and knowledge storage sun2025redeep , respectively..
8 Discussion
Figure 4: The compute-accuracy trade-off on MuSiQue for different metrics and ARC-JSD on 4 LLM backbones with GFLOPs $\mathrm{log}_{10}$ scale per sample.
Comparison JSD with KL, Wasserstein, TV and MMD.
In Fig. 4, we conduct an ablation study among different metrics, which shows that JSD works better than other metrics regarding compute-accuracy trade-off. Here, we discuss some possible reasons: KL divergence will explode whenever ablated run assigns $\approx 0$ probability to a token when full run uses (it is common in deep layers of LLMs). The unbounded scale makes it impossible to compare “how much layer 7 changed” to “how much layer 28 changed”; TV distance is bounded but too coarse, which means that two distributions that swap 5% mass on high-entropy tails give the same TV as two distributions that shift 5% mass off the top-1 token, yet the latter wrecks the answer; Wasserstein needs a distance between tokens. There is no canonical ground metric on a 152K vocabulary (Qwen2-7B-Instruct version), and any choice (e.g., edit distance, embedding cosine, etc.) injects an orthogonal modelling assumption and costs $O(V^{3})$ per layer; MMD always requires a kernel and a feature map to measure a Reproducing kernel Hilbert space (RKHS) norm, which is not tied to likelihood or entropy. It also needs a notion of distance between tokens to build the kernel (See Appendix J for details and examples).
What if all JSD scores are very small?
When all scores are very small, it is the attribution. Small everywhere is not an error, and it means that RAG answers from parametric memory or retrieved passages are irrelevant. In those cases, we prefer to return “no evidence passage was used” rather than force-label the least-bad one. Practically, we can flag the answer with “low-evidence” when all sentence-JSD < 0.02 bits ($\approx$ median noise). The benefit than a threshold is that we can distinguish “no context used” from “weak but present context” without having to guess a universal cut-off. We could use that signal to re-query or warn the user, which is in practice a more faithful and safer behaviour than picking the least-small score.
@@ 9 Conclusion
We introduce ARC-JSD, an inference-time JSD–based metric that directly attributes RAG responses to their source sentences with no fine-tuning or surrogate models needed. Across diverse QA benchmarks and instruction-tuned LLMs, ARC-JSD outperforms different baselines in attribution accuracy while cutting computational overhead; when paired with the Logit Lens, it even pinpoints the specific attention heads and MLPs driving those attributions, advancing the mechanistic interpretability and transparency of RAG systems.
@@ Acknowledgement
Appendix A Appendix
Appendix B Broad Impact
RAG systems underpin a wide range of everyday activities, from itinerary planning and news aggregation to document drafting, by combining LLMs reasoning with evidence retrieved from external sources. Yet, the practical value of these systems hinges on our ability to verify that each generated statement is genuinely grounded in the retrieved material. The proposed post‑hoc ARC-JSD method offers a lightweight, modular solution to this problem. Because ARC‑JSD can be seamlessly integrated into any open‑source RAG pipeline, it provides developers and researchers with an immediate way of auditing attribution fidelity, thereby strengthening the transparency, reliability, and ultimately the public trust in RAG‑based applications.
Appendix C Limitations
Our work focuses on the analysis to (i) identify the context sentences that most strongly influence a RAG model’s output and (ii) attribute that influence to specific attention heads and MLP layers via a JSD-based metric. Two important directions, therefore, remain unexplored. First, our layer‑level view does not reveal which individual neurons within the MLPs mediate context attribution; techniques such as sparse autoencoder (SAE) probing could provide the necessary resolution. Second, we have not yet examined whether our proposed ARC-JSD method as a tool can be generalised to other tasks, such as membership inference attack. This is an exciting direction if the JSD attribution score can be used to locate which data is seen during training.
Appendix D Details of the Internal Mechanisms of LLMs
We consider the standard autoregressive Transformer architecture used in LLMs, originally introduced by [32] and subsequently analysed in a series of mechanistic studies [13, 8, 12, 7, 20, 21, 44]. Given a prompt of length $T$, the input tokens $(t_{1},\dots,t_{T})$ from the context-query pair, each drawn from a vocabulary $\mathcal{V}$, are mapped to dd‑dimensional embedding vectors $\mathbf{x}^{0}_{i}\in\mathbb{R}^{d}$, where the embedding matrix $W_{E}\in\mathbb{R}^{|\mathcal{V}|\times d}$.
@@ Appendix D Details of the Internal Mechanisms of LLMs
where $\mathbf{a}^{\ell}_{i}$ and $\mathbf{m}^{\ell}_{i}$ denote the contributions of the attention and MLP sub‑modules, respectively. [^5] Layer normalisation preceding each sub‑module is omitted here for clarity.
After the final layer, a LayerNorm $\sigma(\cdot)$ and the unembedding matrix $W_{U}\in\mathbb{R}^{d\times|\mathcal{V}|}$ produce the next‑token distribution
@@
A concise per‑head summary is $\mathbf{a}^{\ell,h}_{i}\;=\;\sum_{j}\mathbf{a}^{\ell,h}_{i,j}$.
@@
Appendix E Experimental Details
@@
Appendix F GPT-4.1 as Judge for Comparison between Generated Responses of RAG models and Gold Answers from Datasets
@@
Appendix G Compute-accuracy Trade-off Between Different Baselines and Our ARC-JSD
We mainly compare the compute-accuracy trade-off between different baselines and our proposed ARC-JSD when attributing responses to relevant context. As Figure 5 and 6 show, our ARC-JSD method can achieve up to 3-fold speedup compared to other baselines. In addition, ARC-JSD is consistently Pareto-optimal over different LLM backbone sizes.
Figure 5: The compute-accuracy trade-off on Hotpot QA dataset for 4 baselines and ARC-JSD on 4 LLM backbones
Figure 6: The compute-accuracy trade-off on TyDi QA dataset for 4 baselines and ARC-JSD on 4 LLM backbones
Appendix H Examples of ARC-JSD Context Attribution
@@
The word “mosquito” (formed by mosca and diminutive -ito)[2] is Spanish for “little fly”. [3] Mosquitoes have a slender segmented body, a pair of wings, three pairs of long hair-like legs, feathery antennae, and elongated mouthparts.
@@
His mother’s name was not recorded in history. Title: Mauro Scocco Mauro Scocco (born 11 September 1962) is a Swedish pop artist of Italian descent. He has been described as “one of the sharpest songwriters in Sweden”.
@@
It was written by Bieber, Julia Michaels, Justin Tranter and BloodPop, with production handled by BloodPop. The song was released through GENPOP Corporation, RBMG Records, School Boy Records, Def Jam Recordings and Republic Records on August 17, 2017. It is Bieber’s first single credited as lead artist to be released in over a year, the previous being "Company", which was released in March 2016 for his studio album "Purpose" (2015).
@@
It is also known in Brazil as "Chapolin", "Vermelhinho" ("Little Red One") and "Polegar Vermelho" ("Red Thumb") in allusion to the famous fairy tale character Tom Thumb. Title: Albertsons Albertsons Companies LLC is an American grocery company founded and based in Boise, Idaho. It is privately owned and operated by investors, including Cerberus Capital Management.
Appendix I JSD-based Mechanistic Insights for Located Attention Heads and MLPs
@@
Appendix J Comparisons of JSD with KL, Wasserstein, TV and MMD in Detail
Direct log-probability or KL Divergence.
Most existing baselines, e.g., ContextCite [6], SelfCite [3] and AttriBoT [19], use direct log-probability or KL divergence as metric for context attribution. However, these metrics drop diverges if the masked run assigns $\approx 0$ probability to the token, which is sensitive to highly-skewed token frequencies. Moreover, if JSD is replaced with KL in the Eq. 1 and Eq. 2, it will bring some influence to the attribution impact:
  • Asymmetry / direction choice: We must choose $\text{KL}(P|Q)$ or $\text{KL}(Q|P)$. The ranking of sentences can flip depending on direction. There is no principled reason to prefer one for attribution. Using the symmetrized Jeffreys divergence $\text{KL}(P|Q)+\text{KL}(Q|P)$ removes directionality, but it does not fix the core issues, such as unboundedness, tail sensitivity, numerical instability, and lack of a common scale.
  • Unbounded & numerically unstable: If the ablated run puts (near) zero mass on a token that the full run assigns mass to (that is common at deeper layers), KL explodes or becomes extremely noisy unless we add ad-hoc smoothing. However, this tends to overweight tail events and can produce false positives.
  • Cross-layer incomparability: Because KL or Jeffreys are unbounded, a few positions with tiny denominators dominate the sentence score, i.e., comparing “how much layer 7 changed” vs “layer 28” becomes unstable. JSD’s boundedness is crucial for consistent ranking and aggregation.
Therefore, if we replace JSD with KL, there will be lower precision/recall for “relevant sentence” ranking (it brings more variance, dependence on $\varepsilon$ and direction), which will further lead to low attribution accuracy. It will also tend to disagree more with independent, behaviour-aligned probes (e.g., semantic gain used in our work), although KL divergence has the same FLOPs as JSD.
Wasserstein Distance.
Assume we choose a ground metric $c(a,b)$ over tokens (e.g., token-Hamming, character edit distance, or embedding-cosine cost), and use entropic-regularised Sinkhorn for Wasserstein. When we replace JSD with Wasserstein distance, it will affect attribution:
  • Metric choice drives the result: Edit distance and embedding-cosine will encode orthography or static similarity, not decoding behaviour. They may call a move toward a typo-like token “cheap” and a move toward a semantically correct rival “expensive”, which misaligns with which changes actually flip the output (See more detailed discussion below).
  • Hyperparameters matter: Sinkhorn $\varepsilon$ (regularisation) and number of iterations change the scale and ranking. Different reasonable settings can reorder “relevant sentences”.
  • Context dependence missing: A single cost matrix $c(a,b)$ ignores that token meaning is position- and layer-dependent in a transformer-based LLM, which means that we either accept a mismatch or introduce layer-specific cost matrices (which becomes circular and heavy).
So, rankings become sensitive to modelling choices not tied to the LM’s probability geometry, typically reducing correlation with behaviour (semantic gain) and causal precision in context attribution.
For FLOPs comparison, if we use full support for Wasserstein distance, Sinkhorn per pair costs $O(KV^{2})$ operations (and $O(V^{2})$ memory) for $K$ iterations, where $V\sim 152k$. Instead, if we use top-$k$ support trick, we restrict to top-$k$ tokens of $P$ and $Q$ (say $k\in[100,500]$). Cost becomes $O(Kk^{2})$ per $(layer,r_{j})$, plus top-$k$ selection $O(Vlogk)$. This is still orders of magnitude above JSD in practice and adds hyperparameters $k,K,\varepsilon$.
MMD Metric.
[…]
  • Kernel choice = modelling assumption: We need to make multiple choices: Gaussian or Laplace on which embeddings? What bandwidth? Results (and rankings) will vary with these choices.
  • Units & interpretability: Values depend on kernel scale and there is no direct link to entropy or cross-entropy (which govern decoding). The equal mass moves on tail tokens can dominate if the kernel puts them in “diverse” regions, even though they don’t affect behaviour.
  • Edge case: If we set $k(a,b)=\mathbf{1}[a{=}b]$, MMD reduces to $\ell_{2}$ on probabilities, again misaligned with decoding (uniformly weights all coordinates).
So, replacing JSD with MMD will bring more sensitivity to hyperparameters, and it has weaker correlation with behaviour, and less stable cross-layer comparisons than JSD.
For FLOPs comparison, if we use dense kernel, a naive computation is $O(V^{2})$ per $(\text{layer},r_{j})$ (matrix–vector with $K\in\mathbb{R}^{V\times V}$). Instead, if we use Low-rank/Nyström rank $r$, it will cost $O(rV)$ per pair, but we must tune $r$ and store factors. With $r=256$, this is $\sim\!256\times$ the work of JSD’s $O(V)$ reduction, and quality also depends on $r$. We also need to consider plus kernel selection/bandwidth tuning overhead.
Using edit distance or embedding cosine as metrics for Wasserstein or MMD.
When Wasserstein or MMD uses edit distance or embedding cosine as metrics, it has several limitations:
  1. 1.
    Edit distance (token/character level):
    • Tokenisation mismatch: In subword vocabularies, a single semantic change can span many subwords, and edit distance on token strings becomes an artefact of the tokeniser, not semantics.
    • Semantic blindness: For the example: “Paris” -> “Lyon” (same POS, both cities) and “Paris” -> “Party”. At the token level, any substitution has unit cost, so replacing “Paris” with either “Lyon” or “Party” is equally cheap, despite radically different semantic consequences. With subword tokenisation, the cost becomes tokeniser-dependent. Character-level edit distance differentiates orthography (e.g., “Party” is closer to “Paris” than “Lyon”), which misaligns with factual attribution
    • Decoding irrelevance: The decoder’s choice is driven by probability mass, not string operations. A small edit distance can correspond to a huge shift in probability, and vice versa.
  2. 2.
    Embedding-cosine ground metrics:
    • Context dependence: Token meaning in transformers is contextual. A static vocab-level embedding (or even the unembedding vectors) is not the representation used at the position/layer where attribution is measured. A faithful ground metric would need position- and layer-specific distances, which will explode in complexity and introduce circularity.
    • Anisotropy & polysemy: Cosine distances in high-dimensional language embeddings are known to concentrate and to blur senses, which means that “nearby” vectors can still correspond to different factual claims. Wasserstein might then deem a large semantic change “cheap to move,” underestimating its effect on generation.
    • Tunable choices: Which embedding? Which layer? Do we normalise? Each choice changes the cost matrix and can alter the ranking of “relevant” layers and context sentences, which is exactly the orthogonal modelling assumption we seek to avoid.
TV Metric.
Here, we provide a simple example to explain why TV distance is not an ideal metric to use for context attribution.
The definition of TV distance for two discrete distributions $P,Q$ over the same vocabulary is:
$\mathrm{TV}(P,Q)=\frac{1}{2}\sum_{t}|P(t)-Q(t)|$ (13)
Here, TV measures the total amount of probability mass moved, but not where it moved.
For any decoding methods used in LLMs, they are more affected by the position where probability mass moved, e.g., greedy decoding picks the token with the largest probability, or sampling and beam search are also dominated by how mass is distributed among the top few tokens.
Here is one example to consider a single decoding step with three candidate tokens: $t_{1}$ = the ground-truth/desired token, $t_{2}$ = a strong competitor, $t_{3}$ = a low-probability tail token.
Let the full-context distribution at one decoding step be:
$P=\bigl[p(t_{1}),p(t_{2}),p(t_{3})\bigr]=[0.52,0.43,0.05]$ (14)
Consider two different ablated distributions that both move the same amount of mass $\varepsilon=0.05$:
Case A— move mass in the tail (does not flip the output prediction):
Shift $\varepsilon$ from $t_{3}$ (tail) to $t_{2}$: i.e., $Q_{\text{tail}}=[0.52,0.48,0.00]$. TV calculation will be:
$\mathrm{TV}(P,Q_{\text{tail}})=\tfrac{1}{2}\bigl(|0.52{-}0.52|+|0.43{-}0.48|+|0.05{-}0.00|\bigr)=\tfrac{1}{2}(0+0.05+0.05)=0.05.$ (15)
When we use greedy choice, we still choose $t_{1}$ because 0.52 remains the largest.
Case B — move mass off the top onto its nearest competitor (does flip the output prediction):
Shift the same $\varepsilon=0.05$ from $t_{1}$ to $t_{2}$: $Q_{\text{top}}=[0.47,0.48,0.05]$. TV calculation will be:
$\mathrm{TV}(P,Q_{\text{top}})=\tfrac{1}{2}\bigl(|0.52{-}0.47|+|0.43{-}0.48|+|0.05{-}0.05|\bigr)=\tfrac{1}{2}(0.05+0.05+0)=0.05.$ (16)
When we use greedy choice, it will flip to $t_{2}$ because $0.48>0.47$.
Both perturbations have the same TV = 0.05, but only Case B changes the token the model outputs.
If we move $\varepsilon$ probability from any token i to any token j (and leave all others unchanged), the absolute differences are $|{-}\varepsilon|$ for i, $|{+}\varepsilon|$ for j, and 0 elsewhere, so $\mathrm{TV}(P,Q)=\tfrac{1}{2}\bigl(\varepsilon+\varepsilon\bigr)=\varepsilon$, regardless of which tokens i and j you chose, Which means that TV “sees” only the amount moved, not where it came from or went.
Yet output behaviour depends critically on where the mass moves:
  • The arg-max flips when $p_{j}+\varepsilon>p_{i}-\varepsilon\quad\Longleftrightarrow\quad\varepsilon>\tfrac{1}{2}(p_{i}-p_{j})$. In our numbers, $p_{1}{-}p_{2}=0.09$, so any $\varepsilon>0.045$ flips the token, where Case B does ($\varepsilon=0.05$), Case A does not.
  • For sampling, the log-odds change by $\Delta\log\frac{p_{i}}{p_{j}}=\log\frac{p_{i}-\varepsilon}{p_{j}+\varepsilon}-\log\frac{p_{i}}{p_{j}}$, which is large and negative only when you move mass between the top competitors (Case B), not when you shuffle tail mass (Case A). But TV assigns both moves the same distance.
In addition, we conduct ablation studies to compare different metrics with JSD on compute-accuracy trade-off on different datasets with 4 LLM backbones. As shown in Fig. 13 and Fig. 14, JSD yields a better compute-accuaracy trade-off than other metrics.
Figure 13: The compute-accuracy trade-off on Hotpot QA for different metrics and ARC-JSD on 4 LLM backbones with GFLOPs $\mathrm{log}_{10}$ scale per sample.
Figure 14: The compute-accuracy trade-off on TyDi QA for different metrics and ARC-JSD on 4 LLM backbones with GFLOPs $\mathrm{log}_{10}$ scale per sample.
@@
Appendix K Case Studies of Attention and MLP’s Contribution for Each Response Token
@@
Figure 15: The projection of $\mathbf{x}_{i}^{\ell,\text{mid}}$ and $\mathbf{x}_{i}^{\ell,\text{post}}$ via Logit Lens to vocabulary space from layer 20 to layer 27 of Qwen2-1.5B IT in TyDi QA data sample, where the generated response $\mathcal{R}$ is “The Meiji Restoration took place in Japan.”. Each cell shows the most probable token decoded via Logit Lens.
Figure 16: The projection of $\mathbf{x}_{i}^{\ell,\text{mid}}$ and $\mathbf{x}_{i}^{\ell,\text{post}}$ via Logit Lens to vocabulary space from layer 20 to layer 27 of Qwen2-1.5B IT in TyDi QA data sample, where the generated response $\mathcal{R}$ is “A mosquito has two wings.”. Each cell shows the most probable token decoded via Logit Lens. The colour indicates the probability of the decoded token of the corresponding $\mathbf{x}_{i}^{\ell,\text{mid}}$ or $\mathbf{x}_{i}^{\ell,\text{post}}$.